OpenStudy (anonymous):

Which of the following represents the graph of f(x) = 4x + 2 graph of exponential rising up to the right, through the point 0, negative 3 graph of exponential rising up to the right, through the point 0, 0 graph of exponential rising up to the right, through the point 0, 1 graph of exponential rising up to the right, through the point 0, 3

OpenStudy (candyme):

I often find it easier to graph the equation and compare it to the pictures shown. What would you do first to graph this?

OpenStudy (anonymous):

solve for y

OpenStudy (anonymous):

y=4x+2 would be 4x=y+2?

OpenStudy (candyme):

The equation you typed is not exponential. Did you mean to type it differently?

OpenStudy (anonymous):

idk tbh im not good at these at all

OpenStudy (anonymous):

i just thhout solve for y

OpenStudy (candyme):

\[f(x)=4^{x+2}\] Does your equation look like this or the was it was typed in your question?

OpenStudy (anonymous):

its f(x)=4^x+2 sorry

OpenStudy (candyme):

Okay, that makes more sense. I would start by finding your vertical asymptote. Do you know how to do that?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

set it equl to 0?

OpenStudy (candyme):

You find it by letting x be 0. So substitue 0 in for x

OpenStudy (anonymous):

ok lol one sec

OpenStudy (anonymous):

so it be 2

OpenStudy (anonymous):

hellpp